# Different Inhomogeneous Evolutionary Histories for Uranus and Neptune

### v1

@@ 4 Evolution of Uranus and Neptune
Section 3 demonstrated that initial envelope entropies of above 3.5 kB baryon-1 convectively mix, increasing their internal luminosity significantly compared to models which do not mix. Thus, depending on the rock mass fraction abundance, we chose initial entropy profiles which stabilized the outer composition gradient of Uranus. This is a similar idea to the results of VH20, which preserved an outer composition gradient in a Uranus model to keep the interior heat within the planet and allow the outer regions to cool. In contrast, Neptune’s higher luminosity suggests that its outer composition gradient requires more substantial internal energy replenishment. To reduce the convective stability of Neptune’s outer compositional gradient, we assigned interior entropy profiles higher than Uranus for a particular $Z$ and rock mass fraction distribution. Consequently, a typical Neptune model’s initial temperature profile is approximately 5,000 K hotter than that of Uranus.
@@
Figure 4 showcases the evolution of temperature (top panels) and density (bottom panels) profiles for preferred models of Uranus (left panels) and Neptune (right panels). The selected initial conditions yield a present-day Uranus model consistent with its observed radius, effective temperature, intrinsic thermal flux, and gravity harmonics, as depicted in Figure 5. Due to convective stability imparted by the outer compositional gradient, the outer layers of Uranus (M$>0.92$ MP) cool to temperatures under 1000 K. The interior temperatures stabilize between approximately 6,500 K and 16,000 K and do not cool significantly. This is consistent with the Uranus models of VH20. The top panel of Figure 4 compares these interior temperatures to the phase boundaries of water, ammonia, and methane, showing that Uranus’s interior remains within their supercritical phases. At approximately 0.8 Gyr, Uranus’s temperature profile crosses below the hydrogen-water miscibility curve of Gupta et al. (2024) (blue dotted line), suggesting potential water phase separation. The inferred compact rocky core, characterized by $Z=1$, is approximately 0.5 M. Additionally, the inferred rock mass fraction throughout the inner envelope is 0.35, resulting in an overall ice-to-rock ratio of about 1.5. The inferred moment of inertia factor of the Uranus model is 0.2222, in agreement with the past work of Nettelmann et al. (2013) and VH20.
@@
Conversely, Neptune’s initial entropy profile sufficiently destabilizes the outer compositional gradient, leading to convective mixing at 0.5 Gyr. This mixing triggers a sudden release of interior heat to Neptune’s outer regions. As depicted in the right column of Figure 5, this rapid heat release transiently elevates Neptune’s temperature and luminosity, analogous to the hot model depicted in red in Figure 3. Following this brief episode, the outer regions ($\gtrsim$ 0.4 MP) of the structure cool adiabatically and the radius swiftly contracts, reaching the expected luminosity, effective temperature, and radius at the current age of 4.56 Gyr. At this stage, Neptune’s outer envelope attains temperatures of 10,000 K at 30% of its interior mass and 2,000 K at 60%. In comparison, the corresponding mass coordinates in the Uranus model maintain temperatures of approximately 13,000 K and 8,000 K, respectively. Such internal temperatures in Neptune’s envelope could permit ice phase transitions at later evolutionary stages, as illustrated in Figure 4 (right column). Consequently, portions of Neptune’s envelope might undergo a supercritical-to-ice phase transition. Additionally, Neptune may experience hydrogen-water phase separation early in its evolution, since it cools to temperatures below the hydrogen-water miscibility curve from Gupta et al. (2024). The cooler envelope temperatures at later ages suggest a larger fraction of Neptune’s envelope may be undergoing water depletion. The envelope rock mass fraction in the presented Neptune model is 5% by mass, with an inferred core mass of 3.5 M. The inferred moment of inertia factor of the Neptune model is 0.225.
@@
Figure 5 indicates that the primary evolutionary changes in the Uranus model occur within the initial 0.6 Gyr. The internal flux luminosity of the model aligns closely with the observational data from Pearl & Conrath (1991) and the more recent measurements by Irwin et al. (2025) and Wang et al. (2025). Notably, this model accurately reproduces the refined gravity harmonic measurements reported by Jacobson (2025), which are informed by updated orbital data of Uranus’s rings and moons from French et al. (2024). The updated measurement from Jacobson (2025) represents an order-of-magnitude improvement in precision compared to Jacobson (2014). The Uranus model in Figures 4 and 5 differs by only 0.1% and 0.4% from the observed $J_{2}$ and $J_{4}$ values, respectively. Given the absence of substantial convective mixing in Uranus, its outer heavy element abundance, represented by the dashed line in the top row of Figure 5, remains relatively stable throughout its evolution.
@@
Inferring the density profiles of Uranus and Neptune has been subject of extensive research as a way to constrain the interior structure, mainly by shifting the density profiles to match the gravity harmonic measurements (e.g., Podolak & Reynolds, 1981; Podolak et al., 1991, 1995; Marley et al., 1995; Helled et al., 2011; Movshovitz & Fortney, 2022; Lin et al., 2024). Figure 6 shows the model density profiles of Uranus (top) and Neptune (bottom) at 4.56 Gyr compared with the baseline distribution range proposed by Movshovitz & Fortney (2022) and the density profiles of Nettelmann et al. (2013). The Uranus and Neptune models are consistent with the baseline values of Movshovitz & Fortney (2022). Our Uranus and Neptune models are less dense than the 3-layer models of Nettelmann et al. (2013) due to their higher internal temperatures and ice mixture compositions.
The evolution models in this section are the first inhomogeneous, non-adiabatic models of Uranus and Neptune which incorporate formation-informed initial conditions, including methane-ammonia-water-rock ice mixtures throughout the envelopes of both planets. In addition, these evolution models are the first to be informed by the recent empirical work of Lin et al. (2024), which suggests that the interiors of both planets may include higher amounts of H-He-ice mixtures to better explain the gravity harmonics. The Uranus and Neptune models shown in Figures 4 and 5 are also the first to agree with all the observable parameters in Table 3 simultaneously within reasonable uncertainty. Particularly, the Uranus model is well within range of the updated internal flux measurements of Irwin et al. (2025) and Wang et al. (2025), as well as the updated gravity measurements of Jacobson (2025).
@@ 5 Conclusion
Formation models indicate that both Uranus and Neptune could have initially possessed substantial outer compositional gradients (Valletta & Helled, 2022). To account for Uranus’s low internal luminosity, such a gradient likely restricts heat transfer from the interior, maintaining a cooler exterior (Nettelmann et al., 2016; Vazan & Helled, 2020; Scheibe et al., 2021). Neptune’s higher luminosity could be explained by convective mixing of its outer compositional gradient formed during its early evolution. The evolution of heavy element profiles after formation is strongly influenced by initial thermal or entropy gradients and their spatial distribution (Knierim & Helled, 2024; Tejada Arevalo et al., 2025). Thus, we conclude that the formation models of VH22 can explain the present-day Uranus and Neptune luminosity differences, which could result from variations in their initial entropy/thermal profiles and subsequent heavy element profile evolution. Importantly, we propose that convective mixing could be the primary reason for Neptune’s high internal heat flux.
@@
    In agreement with VH20, the interior regions of Uranus ($\gtrsim 1$ kbar) have likely remained hot and supercritical since formation, provided its outer composition gradient remains convectively stable. The requirement of long-term convective stability is a consequence of 1) the Uranian gravity data (which adiabatic and homogeneous models can not reproduce), 2) the need for initial $Z$ composition gradients and formation models, and 3) its measured internal flux.
@@
The prediction that Neptune may have a higher atmospheric metallicity than Uranus is supported by atmospheric measurements (e.g., Figures 7 and 2 of Guillot et al., 2023; Helled & Howard, 2024, respectively). The atmospheric metallicity measurements of Uranus and Neptune range between $\sim$30–74 and $\sim$40–93 times the solar abundances respectively. [^3] Ranges are based on C/H and S/H ALMA, VLA, HST, and Keck measurements listed in Table 2 of Guillot et al. (2023). In addition to atmospheric measurements, interior models predict that Neptune’s outer metallicity could be higher than that of Uranus. For example, early work by Podolak et al. (1995) predicted that the atmospheric composition of Neptune would be very different than that of Uranus. Guided only by static interior density models and early gravity data, Podolak et al. (1995) could not explain this difference at the time. Another example is the models of Nettelmann et al. (2013), whose Neptune models show a greater outer heavy element abundance than their Uranus models. Superadiabatic atmospheric temperature profiles studied by Leconte et al. (2017) and Cavalié et al. (2017) show that the O/H ratio of Uranus should be less than 160 times the solar value, and about 540 times solar in Neptune.
@@
The geometric albedo spectrum features of Neptune are deeper than that of Uranus at shorter wavelengths (Karkoschka, 1994, 1998). The early work of Karkoschka (1998) found that both Uranus and Neptune show a greater abundance of methane in their atmospheres compared to Jupiter and Saturn, but methane features in Neptune are more pronounced than in the Uranus. Since geometric albedo is a weak function of abundance (Madhusudhan & Burrows, 2012; Burrows, 2014), small differences in spectral depth indicate greater abundance differences. Hence, the differences between Uranus and Neptune at shorter wavelengths imply a greater methane abundance in Neptune’s atmosphere.
@@
We emphasize that the “ices” within Uranus and Neptune may deviate from a strictly solar composition if they experience disassociation and phase separation (Cheng et al., 2021, 2023; Militzer, 2024). Nevertheless, the interior temperatures of these planets may not be sufficiently low to promote the formation of distinct layered structures. Complete phase separation into distinct ice layers likely requires colder temperatures, potentially associated with superionic phase transitions (Millot et al., 2018), as illustrated by the example static 3-layer profiles shown by Militzer (2024), gathered from Redmer et al. (2011), to propose the phase separation of ammonia and methane from water. Past evolution models (Vazan & Helled, 2020) and recent empirical gravitational harmonic analysis (Lin et al., 2024) support the persistence of supercritical ice-rock mixtures at higher temperatures, consistent with the findings of this study. Thus, the interior conditions of Uranus may remain too warm for substantial layering and core freezing, as suggested by Stixrude et al. (2021). However, Figure 4 indicates that supercritical-to-solid ice phase transitions of water, ammonia, and methane could occur in the outer envelope of Neptune.
In this work, we used analytical atmospheric fits from Graboske et al. (1975) and Hubbard (1977) as presented by Guillot et al. (1995). Past work (e.g., Chen et al., 2023; Mankovich & Fortney, 2020) has shown that different atmospheric models significantly affect evolutionary results for Jupiter and Saturn. Self-consistent irradiation models with clouds and up-to-date opacities can make a significant difference ($\gtrsim 6\%$) in the evolution of Jupiter and Saturn (Chen et al., 2023). Particularly, the presence of clouds introduces condensation features which result in deviations from models without clouds (Sudarsky et al., 2000; Chen et al., 2023). Since clouds are expected in the troposphere of both planets (Gautier et al., 1995), clouds could, hence, affect the thermal evolution of Uranus and Neptune. Future research incorporating more sophisticated atmospheric treatments—including solar irradiation, high-metallicity opacities, and the effects of clouds (Chen et al., 2023; Tejada Arevalo et al., 2024; Sur et al., 2025)—are needed. Importantly, better boundary conditions in the presence of convective mixing may slow down the rate of cooling of Neptune instead of abruptly heating up the outer layers. This effect could be more pronounced at low temperatures where hazes and condensates could increase the geometric albedos (Fortney et al., 2011). Evolutionary models with irradiation, metallicity-dependent atmospheric boundary conditions with the effects of clouds will be featured in future work (Tejada Arevalo et al. 2025c, in prep).
@@
Another evolutionary process that affects gas giant planet evolution is H-He phase separation. Instead of H-He separation, water-hydrogen phase separation is expected in both planets (See Figure 4). Future studies will incorporate water and hydrogen phase separation, potentially impacting structural and magnetic evolution (Bailey & Stevenson, 2021; Gupta et al., 2024; Amoros et al., 2024). Hydrogen-water immiscibility, particularly in outer H-He-rich regions at pressures of 10–30 GPa and temperatures below $\sim$6,000 K (Gupta et al., 2024; Amoros et al., 2024), could influence planetary luminosity and composition profile evolution. Notably, water may separate independently from ammonia and methane, which themselves may become immiscible with water at lower temperatures (Militzer, 2024). This work suggests that if hydrogen-water phase separation occurs (or even ice phase separation), then the interior temperatures of Neptune may provide more favorable conditions.
@@
This research was funded by the Center for Matter at Atomic Pressures (CMAP), a National Science Foundation (NSF) Physics Frontier Center under Award PHY-2020249. Any opinions, findings, conclusions, or recommendations expressed herein are those of the authors and do not necessarily reflect NSF views. RTA acknowledges support from the Ford Foundation Predoctoral Fellowship. RTA is grateful to Drs. Adam Burrows, Akash Gupta, Ankan Sur, and Yubo Su for lively discussions, critical reviews and feedback, expertise, and academic mentorship, which significantly influenced this work. RTA is grateful to Dr. Mandy Bethkenhagen for providing equation of state data and valuable discussions. Additionally, RTA thanks Dr. Matthew Coleman and Princeton Research Computing for computational and technical assistance.
@@
Initial Outer Metallicty ($Z_{\rm out}$) 0.05–0.45 0.05–0.25
@@
Figure 6: Present-age density profiles of the Uranus and Neptune models presented in Section 4. The gray lines indicate the baseline range of allowed density profiles predicted by Movshovitz & Fortney (2022, MF2020) based on the gravity data of Jacobson (2014, 2009). Our Uranus and Neptune models are consistent with this range. The red dashed density profiles are those of Nettelmann et al. (2013, N13). Our models are overall less dense than the 3-layer models of N13 due to their higher interior temperatures and ice mixtures.

### v3

@@ 4 Evolution of Uranus and Neptune
Section 3 demonstrated that initial envelope entropies at $\sim$3.8 kB baryon-1 can destabilize the outer composition gradient. [^5] This interior entropy value applies only to the ice mixture ratio being used here. This entropy value is different for different compositions. The goal of this work is not to firmly constrain these numbers since they are subject to all the uncertainties related to Uranus and Neptune. This is a similar idea to the results of VH20, which preserved an outer composition gradient in a Uranus model to keep the interior heat within the planet and allow the outer regions to cool. In contrast, Neptune’s higher luminosity suggests that its outer composition gradient requires more substantial internal energy replenishment. To reduce the convective stability of Neptune’s outer compositional gradient, we assigned interior entropy profiles higher than Uranus for a particular $Z$ and rock mass fraction distribution. Recall from Section 3 that models immediately above the convective mixing threshold satisfy two requirements demanded by the present-day Neptune at once-- a higher present-day intrinsic flux and a smaller present-day radius compared to Uranus. As such, our standard Neptune initial conditions require higher initial entropies than Uranus by $\sim$0.5–1 kB baryon-1 to destabilize the outer region, depending on the rock mass distribution. Values higher than this range can completely destabilize the profile, leading to a homogeneous and adiabatic evolution, which does not yield the expected gravity data of Neptune (and even smaller radii). These entropy constraints lead to a higher initial interior temperature of Neptune of between 2,000–5,000 K relative to initial interior temperatures of Uranus.
@@
All models are evolved to 5 Gyr and their present-age parameters compared with observations (Table 1). These models are scored and ranked according to their fractional errors with respect to the observations, so the models with the least fractional errors are presented here. An example visualization of the relative fractional errors considered here is shown and discussed in Appendix B, Figure 10. Figure 4 showcases the evolution of temperature (top panels) and density (bottom panels) profiles for preferred models of Uranus (left panels) and Neptune (right panels). The selected initial conditions yield a present-day Uranus model consistent with its observed radius, effective temperature, intrinsic thermal flux, and gravity harmonics, as depicted in Figure 5. Due to convective stability imparted by the outer compositional gradient, the outer layers of Uranus (M$>0.92$ MP) cool to temperatures under 1000 K. The interior temperatures stabilize between approximately 6,500 K and 16,000 K and do not cool significantly. This is consistent with the Uranus models of VH20. The top panel of Figure 4 compares these interior temperatures to the phase boundaries of water, ammonia, and methane, showing that Uranus’s interior remains within their supercritical phases. At approximately 0.8 Gyr, Uranus’s temperature profile crosses below the hydrogen-water miscibility curve of Gupta et al. (2024) (blue dotted line), suggesting potential water phase separation. The compact rocky core, characterized by $Z=1$, is approximately 0.5 M. The rock mass fraction throughout the inner envelope is 0.35, resulting in an overall ice-to-rock ratio of about 1.5. The moment of inertia factor of the Uranus model is 0.2222, in agreement with the past work of Nettelmann et al. (2013) and VH20.
@@
Recall from the previous section that higher initial interior entropies systematically yield higher internal fluxes and smaller radii. A Neptune model must show a higher luminosity and smaller radius relative to the Uranus model. As a result, higher initial entropies than Uranus were chosen for the Neptune models (motivated by the models in Figure 3). This higher initial entropy leads to convective mixing at 0.5 Gyr, triggering a sudden release of interior heat to Neptune’s outer regions. As depicted in the right column of Figure 5, this rapid heat release transiently elevates Neptune’s temperature and intrinsic flux, analogous to the hot model depicted in red in Figure 3. Following this brief episode, the outer regions ($\gtrsim$ 0.4 MP) of the structure cool adiabatically and the radius swiftly contracts, reaching the expected intrinsic flux, effective temperature, and radius at the current age of 4.56 Gyr. At this stage, Neptune’s outer envelope attains temperatures of 10,000 K at 30% of its interior mass and 2,000 K at 60%. In comparison, the corresponding mass coordinates in the Uranus model maintain temperatures of approximately 13,000 K and 8,000 K, respectively. Such internal temperatures in Neptune’s envelope could permit ice phase transitions at later evolutionary stages, as illustrated in Figure 4 (right column). Consequently, portions of Neptune’s envelope might undergo a supercritical-to-ice phase transition. Additionally, Neptune may experience hydrogen-water phase separation early in its evolution, since it cools to temperatures below the hydrogen-water miscibility curve from Gupta et al. (2024). The cooler envelope temperatures at later ages suggest a larger fraction of Neptune’s envelope may be undergoing water depletion. The envelope rock mass fraction in the presented Neptune model is 5% by mass, with a core mass of 3.5 M. The moment of inertia factor of the Neptune model is 0.225.
@@
Figure 5 indicates that the primary evolutionary changes in the Uranus model occur within the initial 0.6 Gyr. The internal flux luminosity of the model aligns closely with the observational data from Pearl & Conrath (1991) and the more recent measurements by Irwin et al. (2025) and Wang et al. (2025). Notably, this model accurately reproduces the refined gravity harmonic measurements reported by Jacobson (2025), which are informed by updated orbital data of Uranus’s rings and moons from French et al. (2024). The updated measurement from Jacobson (2025) represents an order-of-magnitude improvement in precision compared to Jacobson (2014). The Uranus model in Figures 4 and 5 differs by 0.1% and 0.4% from the observed $J_{2}$ and $J_{4}$ values, respectively. Readers should note that the measured uncertainties of the Uranian $J_{2}$ are $\sim$0.002%. Evolution models carry more inherent uncertainties than present-day static models (e.g., Movshovitz & Fortney, 2022; Neuenschwander & Helled, 2022; Lin et al., 2024) whose aim is to match the gravitational harmonic data as close as possible. [^6] For instance, evolutionary models of Jupiter (e.g., Vazan et al., 2018; Müller et al., 2020; Tejada Arevalo et al., 2025; Sur et al., 2025) cannot be evaluated against present-day criteria with the same rigor as static interior models (e.g., Wahl et al., 2017; Nettelmann et al., 2021; Militzer et al., 2022; Militzer & Hubbard, 2024), due to the additional complexities involved in modeling planetary evolution. Unless a more complete physical understanding of a planet’s interior is obtained, evolution models will be more inherently uncertain. Given the absence of substantial convective mixing in Uranus, its outer heavy element abundance, represented by the dashed line in the top row of Figure 5, remains relatively stable throughout its evolution.
@@
Inferring the density profiles of Uranus and Neptune has been the subject of extensive research as a way to constrain the interior structure, mainly by varying the density profiles to match the gravity harmonic measurements (e.g., Podolak & Reynolds, 1981; Podolak et al., 1991, 1995; Marley et al., 1995; Helled et al., 2011; Movshovitz & Fortney, 2022; Lin et al., 2024). Movshovitz & Fortney (2022) computed a baseline of density models for Uranus and Neptune based on the available gravitational harmonic data. The outline of their family of density models ranges from $\sim$2–17 g cm-2 (see their Figure 3). The reader should note that the baseline region of Movshovitz & Fortney (2022) represents the envelope of possible density profiles, not the density profiles themselves. The Uranus and Neptune models in Section 4 at 4.56 Gyr are comfortably consistent with their baseline values. Our Uranus and Neptune models are less dense than the 3-layer models of Nettelmann et al. (2013) due to their higher internal temperatures and ice mixture compositions. Figure 6 shows the model density profiles of Uranus (top) and Neptune (bottom) at 4.56 Gyr compared to other models in the literature (Nettelmann et al., 2013; Vazan & Helled, 2020; Neuenschwander & Helled, 2022; Neuenschwander et al., 2024; Morf et al., 2024; Malamud et al., 2024). Of the density profiles shown in Figure 6, only those of Nettelmann et al. (2013) and Vazan & Helled (2020) are evolution models.
The evolution models in this section are the first inhomogeneous, non-adiabatic models of Uranus and Neptune which incorporate formation-informed initial conditions, including methane-ammonia-water-rock ice mixtures throughout the envelopes of both planets. In addition, these evolution models are the first to be informed by the recent empirical work of Lin et al. (2024), which suggests that the interiors of both planets may include higher amounts of H-He-ice mixtures to better explain the gravity harmonics. The Uranus and Neptune models shown in Figures 4 and 5 are also the first to agree with all the observable parameters in Table 3 simultaneously within reasonable uncertainty. Particularly, the Uranus model is well within the range of the updated internal flux measurements of Irwin et al. (2025) and Wang et al. (2025), as well as the updated gravity measurements of Jacobson (2025).
@@ 5 Conclusion
Formation models indicate that both Uranus and Neptune could have initially possessed substantial outer compositional gradients (Valletta & Helled, 2022). To account for Uranus’s low internal luminosity, such a gradient likely restricts heat transfer from the interior, maintaining a cooler exterior (Nettelmann et al., 2016; Vazan & Helled, 2020; Scheibe et al., 2021). Neptune’s higher luminosity could be explained by convective mixing of its outer compositional gradient formed during its early evolution. Past work on inhomogeneous and non-adiabatic gas giant evolution has shown that lower initial entropies in an entropy gradient better preserve stable regions over evolutionary timescales, while stable regions located farther out the structure tend to convectively mix stable regions (Knierim & Helled, 2024; Tejada Arevalo et al., 2025). These two observations have been applied here. Thus, we conclude that the formation models of VH22 can explain the present-day Uranus and Neptune luminosity differences, which could result from variations in their initial entropy/thermal profiles and subsequent heavy element profile evolution. Importantly, we propose that convective mixing could be the primary reason for Neptune’s high internal heat flux whenever the initial interior entropy destabilizes the outer stable region over evolutionary timescales.
@@
    In agreement with VH20, the interior regions of Uranus ($\gtrsim 1$ kbar) have likely remained hot and supercritical since formation, provided its outer composition gradient remains convectively stable. The requirement of long-term convective stability is a consequence of 1) an expected non-adiabatic and inhomogeneous interior (Scheibe et al., 2019; Neuenschwander & Helled, 2022; Neuenschwander et al., 2024; Lin et al., 2024) and 2) the need to explain its low measured internal flux (Nettelmann et al., 2016; Vazan & Helled, 2020).
@@
The prediction that Neptune may have a higher atmospheric metallicity than Uranus is supported by atmospheric measurements (e.g., Figures 7 and 2 of Guillot et al., 2023; Helled & Howard, 2024, respectively). The atmospheric metallicity measurements of Uranus and Neptune range between $\sim$30–74 and $\sim$40–93 times the solar abundances respectively. [^7] Ranges are based on C/H and S/H ALMA, VLA, HST, and Keck measurements listed in Table 2 of Guillot et al. (2023). In addition to atmospheric measurements, interior models predict that Neptune’s outer metallicity could be higher than that of Uranus. For example, early work by Podolak et al. (1995) predicted that the atmospheric composition of Neptune would be very different from that of Uranus. Guided only by static interior density models and early gravity data, Podolak et al. (1995) could not explain this difference at the time. Another example is the models of Nettelmann et al. (2013), whose Neptune models show a greater outer heavy element abundance than their Uranus models. Super-adiabatic atmospheric temperature profiles studied by Leconte et al. (2017) and Cavalié et al. (2017) show that the O/H ratio of Uranus should be less than 160 times the solar value, and about 540 times solar in Neptune.
@@
The geometric albedo spectrum features of Neptune are deeper than that of Uranus at shorter wavelengths (Karkoschka, 1994, 1998). The early work of Karkoschka (1998) found that both Uranus and Neptune show a greater abundance of methane in their atmospheres compared to Jupiter and Saturn, but methane features in Neptune are more pronounced than in Uranus. Since geometric albedo is a weak function of abundance (Madhusudhan & Burrows, 2012; Burrows, 2014), small differences in spectral depth indicate greater abundance differences. Hence, the differences between Uranus and Neptune at shorter wavelengths imply a greater methane abundance in Neptune’s atmosphere.
@@
We emphasize that the “ices” within Uranus and Neptune may deviate from a strictly solar composition if they experience disassociation and phase separation (Cheng et al., 2021, 2023; Militzer, 2024). Nevertheless, the interior temperatures of these planets may not be sufficiently low to promote the formation of distinct layered structures. Complete phase separation into distinct ice layers likely requires colder temperatures, potentially associated with superionic phase transitions (Millot et al., 2018), as illustrated by the example static 3-layer profiles shown by Militzer (2024), gathered from Redmer et al. (2011), to propose the phase separation of ammonia and methane from water. Past evolution models (Vazan & Helled, 2020) and recent empirical gravitational harmonic analysis (Lin et al., 2024) support the persistence of supercritical ice-rock mixtures at higher temperatures, consistent with the findings of this study. Thus, the interior conditions of Uranus may remain too warm for substantial layering and core freezing, as suggested by Stixrude et al. (2021). However, Figure 4 indicates that supercritical-to-solid ice phase transitions of water, ammonia, and methane could occur in the outer envelope of Neptune. We highlight that the specific entropy quantities required to destabilize any given stable layer depend on the scale of the stabilizing entropy gradient (see end of Section 3), which depends on the heavy element equation of state. Models with different volatile abundances relative to water affect the initial conditions and final state of a planet significantly (see Appendix A, Figure 8). However, the general physical idea that a sufficiently high initial interior entropy (i.e., a sufficiently low entropy gradient) convectively mixes an initial $Z$ profile gradient over evolutionary timescales is independent of ice mixture ratios. The Uranus mass models with different ice mixtures in Figure 8 could, in principle, be brought to match the radius and gravity data if initialized at higher entropies, and could convective mix if initialized at higher initial entropies. A future systematic study on the dependence on volatile ratios, rock mass fractions, and other parameters is surely warranted. Therefore, the general main result of this work holds for any mixture of ices and rocks mixed with hydrogen and helium.
A possible scenario which could lead to a higher initial interior entropy for Neptune is a giant impact. Giant impacts during the formation of Neptune are a potential explanation for its hotter initial entropy. Such early collisions could modify the thermal and compositional structures of both Uranus and Neptune (Reinhardt et al., 2020). A sufficiently oblique impact can reproduce Uranus’s extreme obliquity and its prograde, coplanar satellite system (Morbidelli et al., 2012), yet subsequent simulations suggest that an impact alone may not fully account for the present satellite orbits (Rufu & Canup, 2022). The work of Reinhardt et al. (2020) predicts that a giant impact could be enough to tilt Uranus’s spin axis and produce an aligned debris disk. Those authors also conclude that a giant impact could render the Neptunian interior adiabatic. The influence of a giant impact on the interior of Uranus, however, is less clear (Reinhardt et al., 2020). Elsewhere in the Solar System, giant impacts have also been invoked to explain Jupiter’s extended core, but recent hydrodynamical studies indicate that even head-on collisions cannot create the observed fuzzy core (i.e., a stable region) (Sandnes et al., 2024; Meier et al., 2025). A giant impact on Neptune early in its evolution is a compatible hypothesis not only to explain its higher initial entropy, but also its higher mass rocky core.
@@
Another process that affects gas giant planet evolution is H-He phase separation (Stevenson, 1975; Fortney & Hubbard, 2003, 2004; Mankovich et al., 2016; Mankovich & Fortney, 2020; Tejada Arevalo et al., 2025; Sur et al., 2025). Instead of H-He separation, water-hydrogen phase separation is expected in both planets (See Figure 4). Future studies will incorporate water and hydrogen phase separation, potentially impacting structural and magnetic evolution (Bailey & Stevenson, 2021; Gupta et al., 2024; Amoros et al., 2024). Hydrogen-water immiscibility, particularly in outer H-He-rich regions at pressures of 10–30 GPa and temperatures below $\sim$6,000 K (Gupta et al., 2024; Amoros et al., 2024), could influence planetary luminosity and composition profile evolution. Notably, water may separate independently from ammonia and methane, which themselves may become immiscible with water at lower temperatures (Militzer, 2024). This work suggests that if hydrogen-water phase separation occurs (or even ice phase separation), then the interior temperatures of Neptune may provide more favorable conditions.
Formalizing and expanding the parameter search outlined in Section 4 is a promising topic for future work. The space of possible initial conditions is vast, owing to the many combinations of ice–rock mixtures and the stabilizing entropies they require. Here we show that convective mixing can plausibly account for the divergent evolutions of Uranus and Neptune, and that the relevant initial conditions are, in principle, inferrable once their interior compositions are better constrained. A similar strategy has already been applied to gas-giant exoplanets (Knierim & Helled, 2025). A comprehensive inference analysis for Uranus and Neptune is therefore desirable, but its reliability will depend on forthcoming advances in mixture equations of state and in key microphysical inputs such as ice-mixture conductivities and diffusion coefficients.
In this work, we used analytical atmospheric fits from Graboske et al. (1975) and Hubbard (1977) as presented by Guillot et al. (1995). The albedos of both planets are uncertain and updated albedos are necessary to better model the atmospheric cooling. An updated Uranian albedo of 0.349 has been published recently by Irwin et al. (2025), but this difference of albedo (0.3 from Voyager 2; Pearl et al., 1990) leads to negligible differences in Uranus evolution models. Past work (e.g., Mankovich & Fortney, 2020; Chen et al., 2023) has shown that a different atmospheric albedo from 0.3 to 0.5 significantly affects evolutionary results for Jupiter. Self-consistent irradiation models with clouds and up-to-date opacities can make a significant difference ($\gtrsim 6\%$) in the evolution of Jupiter and Saturn (Chen et al., 2023). Particularly, the presence of clouds introduces condensation features which result in deviations from models without clouds (Sudarsky et al., 2000; Chen et al., 2023). Since clouds are expected in the troposphere of both planets (Gautier et al., 1995), clouds could, hence, affect the thermal evolution of Uranus and Neptune. Future research incorporating more sophisticated atmospheric treatments-- including solar irradiation, high-metallicity opacities, and the effects of clouds (Chen et al., 2023; Tejada Arevalo et al., 2024; Sur et al., 2025), is needed. Importantly, better boundary conditions in the presence of convective mixing may slow down the rate of cooling of Neptune instead of abruptly heating up the outer layers. This effect could be more pronounced at low temperatures where hazes and condensates could increase the geometric albedos (Fortney et al., 2011). Evolutionary models with irradiation and metallicity-dependent atmospheric boundary conditions with the effects of clouds will be featured in future work (Tejada Arevalo et al. 2025c, in prep).
Heat conduction governs energy transport across convectively stable layers when semiconvection (double-diffusive convection) is absent. In this study, we combine the water thermal conductivities of French (2019) with the H-He conductivities described in Sur et al. (2024). Thermal conductivity data for ice–rock mixtures are needed and will be essential for refining heat-flux estimates at compositional interfaces. Should semiconvection occur, reliable values for conductivities, Nusselt numbers, and Rayleigh numbers under the conditions expected in Uranus and Neptune are also required. French & Nettelmann (2019) evaluated some of these parameters and concluded that double-diffusive convection is unlikely in either planet. Consequently, composition transport across stable layers is probably purely diffusive, allowing much of each planet’s primordial heat to remain trapped. Our results are consistent with this picture.
@@
We thank the anonymous referee for their thorough reviews that very much sharpened numerous aspects of the manuscript. This research was funded by the Center for Matter at Atomic Pressures (CMAP), a National Science Foundation (NSF) Physics Frontier Center under Award PHY-2020249. Any opinions, findings, conclusions, or recommendations expressed herein are those of the authors and do not necessarily reflect NSF views. RTA acknowledges support from the Ford Foundation Predoctoral Fellowship. RTA is grateful to Drs. Adam Burrows, Akash Gupta, Ankan Sur, and Yubo Su for lively discussions, critical reviews and feedback, expertise, and academic mentorship, which significantly influenced this work. RTA is grateful to Dr. Mandy Bethkenhagen for providing equation of state data and valuable discussions. Additionally, RTA thanks Dr. Matthew Coleman and Princeton Research Computing for computational and technical assistance.
@@
Initial Outer Metallicity ($Z_{\rm out}$) 0.05–0.45 0.05–0.25
@@
Figure 6: Present-age density profiles of the Uranus and Neptune models (thick black lines) presented in Section 4 compared with models from the literature (color lines; Nettelmann et al., 2013; Vazan & Helled, 2020; Neuenschwander & Helled, 2022; Neuenschwander et al., 2024; Malamud et al., 2024). Of these, only the Uranus models of Nettelmann et al. (2013) and Vazan & Helled (2020) are the result of evolutionary calculations. Others are static inferences. Since these authors produce multiple models with varying of water/rock compositions, we indicate which are plotted in parentheses. Our Uranus models are consistent with literature results. There are fewer Neptune density profiles in the literature, but ours is consistent with the Neptune model of Nettelmann et al. (2013). Neuenschwander & Helled (2022) produced their models using so-called “empirical” EOSes, while Malamud et al. (2024) showed only an example profile instead of a preferred model. The red dashed density profiles are those of Nettelmann et al. (2013, N13). Our models are overall less dense than the 3-layer models of N13 due to their higher interior temperatures and ice mixtures.